OpenStudy (anonymous):

Find the standard form for the equation of a circle (x-h)^2 + (y-k)^2 = r^2 with a diameter that has endpoints (-1, 4) and (5, 0). i just need to find r

OpenStudy (anonymous):

if you find the length of a diameter and divide it by 2 .you will find it's radius .

OpenStudy (anonymous):

well i got (2,2) for h,k do i just plug in the equation to find r?

OpenStudy (anonymous):

well.you only want to find r so you should find the length of a diameter with using this formula: \[D=\sqrt{(y _{2}-y _{1})^2+(x _{2}-x _{1})^2}\]

OpenStudy (anonymous):

so i got sqrt 37

OpenStudy (anonymous):

are you sure? (0-4)^2+(5-(-1))^2=16+36=52 D=sqrt52=sqrt4*13=2*(sqrt13)

OpenStudy (anonymous):

opps i switched them